# Adapting Conformal Prediction to Distribution Shifts Without Labels

### v1

@@ 4 Our methods
In this section, we first propose a method called ECP (Entropy scaled Conformal Prediction), which improves the coverage rate of CP by enlarging its prediction sets using the uncertainty evaluation of the base model itself. Crucially, this notion of uncertainty can be directly minimized and refined through unsupervised TTA, leading to an improved method called EACP (Entropy base-Adapted Conformal Prediction). It is able to both i) recover the desired error rate on many challenging distribution-shifted datasets, while ii) significantly reduce inflated set sizes under increased uncertainty.
@@ 4.1 Scaling conformal scores by uncertainty
Let us start by motivating our idea. A common criticism of standard SplitCP (although typically studied in the regression setting rather than classification) is that the size of the CP prediction set is invariant to the covariate $x$. To provide greater “local adaptivity” (w.r.t. $x$), uncertainty-aware techniques have been proposed to adjust (or scale) the score function $s(x,y)$ based on some notion of uncertainty (or difficulty) that the base model decides at each $x$ (Papadopoulos et al., 2008; Johansson et al., 2015; Lei et al., 2018; Izbicki et al., 2019; Romano et al., 2019; Seedat et al., 2023; Rossellini et al., 2024). This has the effect of inflating the prediction set on the base model’s uncertain regions, and has been shown to improve the conditional coverage rate of CP (Angelopoulos and Bates, 2023; Tibshirani, 2023).
Figure 1: Entropy vs. the softmax score of true label, averaged on each dataset.
@@
Building on these results, our key idea is to apply an analogous uncertainty scaling on the score function, in order to improve the performance of CP under distribution shifts. However, instead of using the uncertainty of the base model at each covariate $x$, we crucially evaluate its uncertainty on the whole distribution-shifted test dataset, effectively aggregating its “localized” uncertainty at each $x$. In other words, instead of aiming for “local adaptivity” (w.r.t. $x$) as in prior works, we aim for the adaptivity w.r.t. the unknown distribution shift.
@@
More concretely, which uncertainty measure do we use on the base model? As discussed above, the ideal dataset-specific uncertainty measure would follow from a “localized” uncertainty measure at each $x$, and a particularly useful one for the latter is the entropy of the model’s predictions,
@@
Previous works have established the relation between such an entropy notion (of the base model) and the magnitude of the distribution shift, showing that larger shifts are strongly correlated with higher entropy (thus uncertainty in the base model) (Wang et al., 2021). We provide a consistent but unique observation in Figure 1, which plots the relation between the entropy (averaged over all $x$ values in the dataset) and the softmax score of the true label (also averaged over $x$), across a range of different datasets. [^4] Darker shades represent greater severity levels of ImageNet-C corruptions; see Section 5 for more details on the datasets. For the true label to be included in the CP prediction set, which is eventually what we aim for, its softmax value should be greater than the CP threshold $\tau_{D}$. From Figure 1, one could see that an increase in entropy is associated with a decrease in the softmax score of the true label, which crucially means that the prediction sets should be larger when the entropy is higher.
Figure 2: The targeted coverage rate $1-\alpha$ vs. the empirical coverage rate, induced by ECP with different $\beta$ values. It shows that simply setting $\beta=1-\alpha$ in ECP consistently works well for all but the most severe distribution shifts (e.g., ImageNet-R and ImageNet-C Contrast Severity 5). Such an observation also holds across various $\alpha$ values, suggesting the general effectiveness of this hyperparameter choice.
Now consider going from the “localized” uncertainty measure $h(x)$ to an uncertainty measure on the test dataset, denoted as $u_{D_{\mathrm{test}}}$. We introduce another design choice: defining $u_{D_{\mathrm{test}}}$ as a specified, $\beta$-quantile of $h(x)$ on the test dataset. Surprisingly, we find that simply setting $\beta$ to the desired coverage rate $1-\alpha$ is a fairly reliable choice in practice (see Figure 2), which gives a robust (over)-estimate of typical $h(x)$ values on the test dataset. We perform all the experiments with this direct relationship to avoid excessive hyperparameter tuning, but it can be further refined if desired.
@@
Now we are ready to use $u_{D_{\mathrm{test}}}$ above to scale the score functions on the test dataset, without label access. The resulting method is named as ECP (Entropy scaled Conformal Prediction).
Formally, let us denote $q_{\beta}(\cdot)$ as the $\beta$-th quantile of its argument, and $h_{D_{\mathrm{test}}}=\{h(x_{1}),...,h(x)_{N_{\mathrm{test}}}\}$ as the entropy values on the test dataset. Then, we compute the “entropy quantile”
$u_{D_{\mathrm{test}}}=q_{1-\alpha}\left(h_{D_{\mathrm{test}}}\right),$ (3)
and on any test sample, modified from Eq.(2), we scale the score function by $\max(1,u_{D_{\mathrm{test}}})$ to form the CP prediction set
$\mathcal{C}_{D_{\mathrm{test}}}\left(x;\tau_{D}\right):=\left\{y^{\prime}\in[k]:s(x,y^{\prime})\cdot\max(1,u_{D_{\mathrm{test}}})\geq\tau_{D}\right\}.$ (4)
Here, we take a maximum with $1$ to ensure that the prediction sets of ECP cannot be smaller than those of standard SplitCP.
The intuition is that a larger distribution shift will result in larger entropy, and thus lead to a correspondingly larger up-scaling of the score function. In this way, more labels have score larger than the fixed CP threshold $\tau_{D}$, and the prediction set grows. This can help mitigate the under-coverage issue of standard SplitCP under distribution shifts (without accessing test labels), and further details are provided in our experiments (Section 5).
@@ 4.2 Optimizing uncertainty using TTA
While our ECP method already improves the coverage of SplitCP on several datasets, it inevitably leads to larger set sizes and, like all post-hoc CP methods, still relies on a fixed base model. To remedy this, we refine ECP using techniques from unsupervised TTA, which are able to update the base model’s parameters on the unlabeled test dataset. While these techniques have been often investigated in the context of top-1 accuracy, we take a different perspective and study their ability to improve set-valued classifiers like conformal prediction.
@@
Concretely, consider the following loss function (of the base model’s parameter $\theta$),
@@
Our insight is that by directly minimizing this loss (through updating $\theta$), we implicitly minimize the entropy quantile $u_{D_{\mathrm{test}}}$ used to scale the score functions. Consequently, the sizes of ECP prediction sets can be reduced, and the coverage rate can be increased.
@@
A number of specific TTA methods have been developed to minimize Eq.(5), while ensuring certain notions of stability. In this work, we leverage the recent ETA (Efficient Test-time Adaptation) method (Niu et al., 2022) as it is a simple approach shown to be effective even in continuously shifting distributions (Press et al., 2023). Adopting this with ECP results in a new CP method, which we call EACP (Entropy base-Adapted Conformal Prediction).
In practice, one could simply call ETA as a subroutine, so we only present its high level idea for the sake of completeness. On a batch of new test data (i.e., $x_{i}$ with a collection of indices $i$), ETA filters the base model’s outputs (i.e., softmax scores) $s(x_{i},\cdot)$ by excluding the outputs similar to those already seen. Then, it reweights the remaining outputs based the associated entropy $h(x_{i})$, with lower entropy (less uncertain) indices receiving higher weights. This leads to a weighted variant of the loss function Eq.(5), which is then minimized by performing a single gradient update. Subsequently, the updated base model is applied to ECP to form the prediction sets of EACP, according to Eq.(4).
In Section 5, we demonstrate that EACP can further improve the performance of ECP, by increasing the coverage rate while maintaining informative set sizes.
@@ 4.3 Uncertainty scaling function
Previously, in Eq.(4) we scale the score functions linearly by the entropy quantile $u_{D_{\mathrm{test}}}$ of the base model. However, in principle this can be adjusted by any (potentially non-linear) function $f(\cdot)$ of the entropy quantile, $f(u_{D_{\mathrm{test}}})$. The best choice of $f(\cdot)$ depends on the unknown relation between $u_{D_{\mathrm{test}}}$ and the $(1-\alpha)$-quantile of the true labels’ conformal scores, denoted as
$\tau_{D_{\mathrm{test}}}:=q_{1-\alpha}\left[s(x,y);(x,y)\in D_{\mathrm{test}}\right].$
Specifically, we aim to find an $f(\cdot)$ satisfying $f(u_{D_{\mathrm{test}}})=\tau_{D}/\tau_{D_{\mathrm{test}}}$.
Figure 3: $u_{D_{\mathrm{test}}}$ versus $\tau_{D}/\tau_{D_{\mathrm{test}}}$ on a log-log scale. On mild and moderate distribution shifts, the linear fit on the log-log plot has slope between 1 and 2. This suggests the effectiveness of using a linear or quadratic function as $f(\cdot)$, which acts on the entropy quantile. However, we also observe that a higher-order polynomial is required on more difficult shifts, such as ImageNet-R.
While finding this optimal $f(\cdot)$ is obviously infeasible without observing the labels in $D_{\mathrm{test}}$, in Figure 3 we empirically evaluate the ideal choice in a post-hoc manner, across different datasets. Recall that both $u_{D_{\mathrm{test}}}$ and $\tau_{D_{\mathrm{test}}}$ depend on the desired error rate $\alpha$. Therefore, for each dataset, we vary $\alpha$ and plot the resulting $u_{D_{\mathrm{test}}}$ versus $\tau_{D}/\tau_{D_{\mathrm{test}}}$ on a log-log scale. Furthermore, if we mildly restrict $f(\cdot)$ to polynomials, then its optimal order can be approximated by the slope of a linear fit on the log-log plot. We only use not-too-small $\alpha$ values for the linear fit, since it is closer to the practice of CP and less prone to noise.
Figure 3 shows that the optimal polynomial order generally increases with the severity of the distribution shift. A larger polynomial order would lead to larger prediction sets. While end-users can refine $f(\cdot)$ based on a preference towards ensuring coverage or small set sizes, we empirically validate that both linear ($\texttt{ECP}_{1}$ / $\texttt{E{\small A}CP}_{1}$) and quadratic ($\texttt{ECP}_{2}$ / $\texttt{E{\small A}CP}_{2}$) scaling are sufficient for coverage and moderate set sizes, on a diverse range of distribution shifts.
@@ 5 Experiments
We conduct experiments across a number of large-scale datasets and neural network architectures. Our setup builds on the standard SplitCP procedure introduced in Section 3, which relies on a held-out, in-distribution, “development set” for calibrating the CP threshold. On ImageNet variants, we split the original ImageNet development set, into a calibration and validation set, resulting in 25,000 calibration points. we refer to Appendix A.1 for more details on the calibration splits and sizes for all datasets.
The conformal threshold is found on the calibration set, and used for all subsequent distribution-shifted datasets. After the conformal threshold is estimated in-distribution, all subsequent steps are unsupervised. We show results on both stationary and continuously shifting distributions.
@@
We compare our proposed ECP and EACP methods to a number of possible baselines. This includes using a pre-calibrated threshold on new distribution-shifted data as in the standard SplitCP approach, using this calibrated threshold while updating the base model using ETA, and an approach we call NAIVE which forms prediction sets by including classes until their cumulative probability is greater or equal to the desired $1-\alpha$ coverage level, without any prior calibration. In all experiments, the target coverage is 0.90. We also analyze linear and quadratic scaling of the uncertainty adjustment, as described in Section 4.3.
@@
While the previous datasets present a single distribution shift, the ImageNet-C (Hendrycks and Dietterich, 2019) dataset allows us to investigate shifts across many types and severities. Specifically, ImageNet-C applies 19 visual corruptions to the ImageNet validation set across four corruption categories — noise, blur, weather, and digital, with five severity levels for each corruption. Please refer to Appendix A.1 for more details on the studied datasets.
5.1 Stationary shifts
Table 1: ECP and EACP can achieve the desired coverage on a number of distribution-shifted datasets, across a variety of imaging domains (ecological, cellular, satellite, etc.). The coverage gap is significantly reduced even on extremely challenging datasets. All results are from ResNet-50 models except FMOW, which uses a DenseNet-121 (Huang et al., 2016). Quadratic uncertainty scaling provides better coverage rates, however, linear scaling results in smaller set sizes.
Coverage Set Size
Dataset SplitCP NAIVE ETA $\texttt{ECP}_{1}$ $\texttt{ECP}_{2}$ $\texttt{E{\small A}CP}_{1}$ $\texttt{E{\small A}CP}_{2}$ SplitCP NAIVE ETA $\texttt{ECP}_{1}$ $\texttt{ECP}_{2}$ $\texttt{E{\small A}CP}_{1}$ $\texttt{E{\small A}CP}_{2}$
ImageNet-V2 0.81 0.88 0.81 0.86 0.91 0.86 0.91 2.5 11.7 2.5 4.2 7.6 4.5 8.7
ImageNet-R 0.50 0.69 0.62 0.61 0.72 0.71 0.80 3.4 20.9 3.0 9.1 23.3 6.8 16.1
ImageNet-A 0.03 0.14 0.05 0.10 0.27 0.14 0.30 3.4 12.7 3.6 7.4 15.1 8.7 10.1
iWildCam 0.84 0.76 0.84 0.84 0.88 0.84 0.89 3.9 2.5 3.8 3.8 5.5 3.7 5.6
RXRX1 0.84 0.48 0.87 0.87 0.90 0.90 0.93 81.8 6.4 100 105 137 133 177
FMOW 0.87 0.83 0.87 0.93 0.96 0.93 0.94 6.2 5.8 6.5 10.3 15.3 11.1 16.4
@@ 5.1 Stationary shifts
Table 1 summarizes our performance on various natural distribution-shifted datasets. We observe that ETA alone is unable to fully recover the desired coverage, however, ECP closes the gap significantly. Coverage is further improved by adaptation through EACP, which helps to reduce set sizes on some datasets – it recovers the desired coverage on all but the most challenging datasets. Here, we also see the trade-off between linear and quadratic uncertainty scaling. We observe $\texttt{E{\small A}CP}_{2}$ consistently achieves higher coverage rates, however this also leads to “over-coverage” on some datasets and thus larger sets. In contrast, $\texttt{E{\small A}CP}_{1}$ leads to lower coverage but also smaller set sizes. This trade-off can be selected by end-users depending on their preference for more accurate or more efficient prediction sets. In subsequent experiments, we will focus on demonstrating the benefit of $\texttt{E{\small A}CP}_{2}$ on coverage, while noting that the observed set sizes are nonetheless practically useful and far from trivial.
Table 2: Coverage on four different corruption types representing each ImageNet-C category. By explicitly leveraging uncertainty under distribution shift, our $\texttt{ECP}_{2}$ method closes the coverage gap on most severity levels. $\texttt{E{\small A}CP}_{2}$ improves on this by refining the uncertainty at test-time, resulting in a recovered coverage rate of 0.90 on nearly all corruption types and severities.
Method Contrast Brightness Gaussian Noise Motion Blur
1 2 3 4 5 1 2 3 4 5 1 2 3 4 5 1 2 3 4 5
NAIVE 0.91 0.89 0.87 0.83 0.76 0.92 0.92 0.91 0.91 0.90 0.88 0.85 0.79 0.69 0.79 0.91 0.90 0.85 0.77 0.71
SplitCP (Sadinle et al., 2019) 0.83 0.78 0.66 0.36 0.09 0.88 0.87 0.86 0.83 0.78 0.79 0.69 0.50 0.26 0.07 0.83 0.74 0.57 0.37 0.27
ETA (Niu et al., 2022) 0.87 0.86 0.84 0.79 0.63 0.88 0.88 0.87 0.86 0.84 0.86 0.82 0.76 0.69 0.54 0.86 0.84 0.80 0.73 0.68
$\texttt{ECP}_{2}$ (ours) 0.93 0.92 0.89 0.79 0.60 0.94 0.94 0.94 0.93 0.92 0.92 0.88 0.80 0.86 0.38 0.94 0.92 0.86 0.75 0.68
$\texttt{E{\small A}CP}_{2}$ (ours) 0.93 0.93 0.93 0.92 0.87 0.93 0.93 0.93 0.93 0.93 0.93 0.93 0.92 0.90 0.84 0.93 0.92 0.92 0.91 0.89
@@
In Table 2, we show fine-grained results of our method’s performance on one corruption type for each ImageNet-C category and across each severity level. Here, we can see the benefit of leveraging an uncertainty notion that can be directly minimized and refined on new test samples. By both adjusting the conformal scores by an entropy-based notion of uncertainty while simultaneously refining this uncertainty, EACP is able to recover the target coverage rate on most corruption types and severities.

### v2

@@ 3 Our methods
In this section, we first propose a method called ECP (Entropy scaled Conformal Prediction), which improves the coverage rate of CP by enlarging its prediction sets using the uncertainty evaluation of the base model itself. Crucially, this notion of uncertainty can be directly minimized and refined through unsupervised TTA, leading to an improved method called EACP (Entropy base-Adapted Conformal Prediction). The latter is able to both recover the desired error rate on many challenging distribution-shifted datasets, and significantly reduce inflated set sizes under increased uncertainty.
@@ 3.1 Scaling conformal scores by uncertainty
Let us start with a high-level motivation. Within the SplitCP framework, an important design objective is local adaptivity: the size of the prediction set $\mathcal{C}(x)$ needs to vary appropriately with the covariate $x$. To this end, standard practice is to adjust the score function $s(x,y)$ based on some notion of uncertainty (or difficulty) that the base model decides at each $x$ (Papadopoulos et al., 2008; Johansson et al., 2015; Lei et al., 2018; Izbicki et al., 2020; Romano et al., 2019; Seedat et al., 2023; Rossellini et al., 2024). This has the effect of inflating the prediction set on the base model’s uncertain regions, and has been shown to improve the more informative conditional coverage rate of CP (Angelopoulos and Bates, 2023; Tibshirani, 2023).
Figure 1: Entropy vs. the softmax score of the true label, averaged on each dataset. Different colors represent different datasets, and darker shades represent greater severity levels of ImageNet-C corruptions. See Section 4 for more details on the datasets.
@@
Inspired by these results, our key idea is to apply an analogous uncertainty scaling on the score function, to improve the performance of CP under distribution shifts. However, instead of using the uncertainty of the base model at each covariate $x$, we draw a crucial connection to unsupervised TTA, and evaluate the base model’s uncertainty on the whole distribution-shifted test dataset $D_{\mathrm{test}}$ – this effectively aggregates its “localized” uncertainty at the test covariates $\{x_{i}\}_{i\in[N]}$. In other words, instead of aiming for “local adaptivity” as in prior works, we use uncertainty scaling to achieve the adaptivity w.r.t. the unknown distribution shift.
@@
More concretely, which uncertainty measure should we use on the base model? As discussed above, the ideal dataset-specific uncertainty measure would follow from a “localized” uncertainty measure at each covariate $x$, and in the context of classification, a particularly useful one is the entropy of the base model’s probabilistic prediction,
@@
Previous works have established the relation between such an entropy notion and the magnitude of the distribution shift, showing that larger shifts are strongly correlated with higher entropy (thus higher uncertainty in the base model) (Wang et al., 2021; Kang et al., 2024). We provide a consistent but unique observation in Figure 1, which plots the relation between the entropy (averaged over all $x$ values in the dataset) and the softmax score of the true label (also averaged over $x$), evaluated on a ResNet-50 model [^3] We fix the base model to ResNet-50 in most of our experiments, unless otherwise specified. and across a range of datasets. For the true label to be included in the CP prediction set, which is eventually what we aim for, its softmax score should be greater than the CP threshold $\tau_{D}$. Figure 1 shows that an increase in entropy is associated with a decrease in the softmax score of the true label, which crucially means that we need to scale up the score function in order to still cover the true label.
Figure 2: The targeted coverage rate $1-\alpha$ vs. the empirical coverage rate, induced by ECP with different $\beta$ values (represented by different colors). It shows that simply setting $\beta=1-\alpha$ in ECP (i.e., the blue dots) consistently works well for all but the most severe distribution shifts (e.g., ImageNet-R and ImageNet-C Contrast Severity 5). Such an observation also holds across various $\alpha$ values, suggesting the effectiveness of this hyperparameter choice.
Now consider going from the “localized” uncertainty measure $h(x)$ to an uncertainty measure on the test dataset $D_{\mathrm{test}}$, denoted as $u_{\mathrm{test}}$. One could use the average $N^{-1}\sum_{i=1}^{N}h(x_{i})$, but to increase the robustness, we define $u_{\mathrm{test}}$ as the $\beta$-quantile of $\{h(x_{i})\}_{i\in[N]}$, where $\beta$ is a hyperparameter. Quite surprisingly, we find that simply setting $\beta$ to the desired coverage rate $1-\alpha$ is a fairly reliable choice in practice (see Figure 2), which gives a robust (over)-estimate of typical $h(x)$ values on the test dataset. We perform all the experiments with this direct relationship to avoid excessive hyperparameter tuning, but it can be further refined if desired.
@@
Now we are ready to use $u_{\mathrm{test}}$ above to scale the score functions on the test dataset, without label access. The resulting method is named as ECP (Entropy scaled Conformal Prediction).
Formally, define $q_{\beta}(\cdot)$ as the $\beta$-th quantile of its argument, and let the base model’s uncertainty measure $u_{\mathrm{test}}$ be the “entropy quantile”
$u_{\mathrm{test}}=q_{1-\alpha}(\{h(x_{i})\}_{i\in[N]}).$ (3)
On any test covariate $x_{i}$, modified from Eq.(2), we scale the score function by $\max(1,u_{\mathrm{test}})$ to form the prediction set
$\mathcal{C}(x_{i}):=\{y\in[k]:s(x_{i},y)\cdot\max(1,u_{\mathrm{test}})\geq\tau_{D}\}.$ (4)
Here, we take a maximum with $1$ to ensure that the prediction sets of ECP cannot be smaller than those of standard SplitCP. The pseudocode is presented as Algorithm 1 in the next subsection.
To recap, the intuition of ECP is that a larger distribution shift will result in larger entropy predicted by the base model, which then leads to a correspondingly larger up-scaling of the score function. In this way, more candidate labels have scores larger than the fixed CP threshold $\tau_{D}$, and the prediction set grows. Without any access to the test labels, this can help mitigate the under-coverage issue of standard SplitCP under distribution shifts, and further details are provided in our experiments (Section 4).
@@ 3.2 Optimizing uncertainty using TTA
While ECP already improves the coverage rate of SplitCP on several datasets, it inevitably leads to larger set sizes and, like typical post-hoc CP methods, still relies on a fixed base model. To remedy this, we refine ECP using entropy minimization (Grandvalet and Bengio, 2004; Wang et al., 2021), a classical idea in unsupervised TTA which updates the base model itself on the unlabeled test dataset. Although such techniques in unsupervised TTA have been investigated in the context of top-1 accuracy, we take a different perspective and study their ability to improve set-valued classifiers like conformal predictors.
@@
Concretely, we first rewrite the entropy $h(x)$ as a loss function w.r.t. the base model’s parameter $\theta$,
@@
Our main idea is to update the base model by minimizing this loss function (or a suitable variant) on the test dataset $D_{\mathrm{test}}$, before applying ECP. This brings two benefits.
  • The updated base model is better suited for the shifted distribution $\mathcal{D}_{\mathrm{test}}$, which generally improves the quality of the prediction sets built on top of it.
  • The base model’s entropy determines the amount of prediction set inflation due to ECP. By directly minimizing the entropy, the resulting prediction sets can be smaller.
@@
A number of specific TTA methods have been developed to minimize entropy, while ensuring certain notions of stability. In this work, we leverage a recent method called ETA (Efficient Test-time Adaptation; Niu et al. 2022), due to its simplicity and effectiveness even under continual distribution shifts (Press et al., 2023). Combining this with ECP results in a new CP method, which we call EACP (Entropy base-Adapted Conformal Prediction).
In practice, one could simply call ETA as a subroutine, so here we only present its high level idea for completeness. First, the test dataset $D_{\mathrm{test}}$ is divided into a collection of batches. On each batch (i.e., $x_{i}$ with a collection of indices $i$), ETA filters the base model’s outputs (i.e., softmax scores) $s(x_{i},\cdot)$ by excluding the outputs similar to those already seen. Then, it reweighs the remaining indices based the associated entropy $h(x_{i})$, with lower entropy (less uncertain) indices receiving higher weights. This leads to a weighted batch variant of the loss function Eq.(5), which is then minimized by performing a single gradient update. Subsequently, the updated base model is applied to ECP to form the prediction sets of EACP, according to Eq.(4).
The combined pseudo-code of ECP and EACP is provided in Algorithm 1. Here we include an uncertainty scaling function $f$ as a small generalization, which acts on the entropy quantile before generating the prediction sets. So far we have only considered the trivial scaling $f(x)=x$, but more choices will be studied in the next subsection.
Algorithm 1 Combined pseudocode of ECP and EACP
Test dataset $D_{\mathrm{test}}=\{x_{i}\}_{i\in[N]}$; trained model with parameter $\theta$ and softmax score $\pi_{\theta}(x)_{y}$; targeted error rate $\alpha$; score threshold $\tau_{D}$ for the error rate $\alpha$, calculated on a calibration dataset $D$; uncertainty scaling function $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$.
if EACP then
     $\theta\leftarrow\texttt{ETA}(\theta,D_{\mathrm{test}})$triangleright TTA sub-routine
end if
$u_{\mathrm{test}}\leftarrow q_{1-\alpha}(\{h(x_{i})\}_{i\in[N]})$triangleright Eq.(3)
$u_{\mathrm{test}}\leftarrow f(u_{\mathrm{test}})$triangleright Uncertainty scaling
for $x_{i}\in D_{\mathrm{test}}$ do
     return $\mathcal{C}(x_{i}):=\{y\in[k]:s(x_{i},y)\cdot\max(1,u_{\mathrm{test}})\geq\tau_{D}\}$triangleright Predict the label set, Eq.(4)
end for
In Section 4, we demonstrate that EACP can further improve the empirical performance of ECP, by increasing the coverage rate while maintaining informative set sizes.
@@ 3.3 Uncertainty scaling function
Figure 3: $u_{\mathrm{test}}$ versus $\tau_{D}/\tau_{\mathrm{test}}$ on a log-log scale. On mild and moderate distribution shifts, the linear fit on the log-log plot has slope between 1 and 2. This suggests the effectiveness of using a linear or quadratic function as $f(\cdot)$, which acts on the entropy quantile. However, we also observe that a higher-order polynomial is required on more difficult shifts, such as ImageNet-R.
In Eq.(4), we essentially scale the score functions linearly by the entropy quantile $u_{\mathrm{test}}$ of the base model. However, this can be adjusted more generally by any (potentially non-linear) function $f(\cdot)$. The best choice of $f(\cdot)$ depends on the unknown relation between $u_{\mathrm{test}}$ and the $(1-\alpha)$-quantile of the ground truth labels’ conformal scores, denoted as [^4] Recall that the data we face at test time is denoted as $\{x_{i},y_{i}\}_{i\in[N]}$, with the observed part (covariates) denoted as $D_{\mathrm{test}}=\{x_{i}\}_{i\in[N]}$.
$\tau_{\mathrm{test}}:=q_{1-\alpha}\left[s(x_{i},y_{i});x_{i}\in D_{\mathrm{test}}\right].$
The optimal $f(\cdot)$ should satisfy $f(u_{\mathrm{test}})=\tau_{D}/\tau_{\mathrm{test}}$.
While finding this optimal $f(\cdot)$ is obviously infeasible without observing the ground truth labels at test time, in Figure 3 we empirically evaluate the ideal choice in a post-hoc manner, across different datasets, in order to demonstrate the insights. Recall that both $u_{\mathrm{test}}$ and $\tau_{\mathrm{test}}$ depend on the desired error rate $\alpha$. Therefore, for each dataset, we vary $\alpha$ and plot the resulting $u_{\mathrm{test}}$ versus $\tau_{D}/\tau_{\mathrm{test}}$ on a log-log scale. If we mildly restrict $f(\cdot)$ to the family of polynomials, then its optimal order can be approximated by the slope of a linear fit on the log-log plot. We only use not-too-small $\alpha$ values (i.e., the lower left corner on the plot) for the linear fit, since it is closer to the typical practice and less prone to noise.
Figure 3 shows that the optimal polynomial order generally increases with the severity of the distribution shift, which is consistent with the fact that a larger polynomial order would lead to larger prediction sets using our methods. While end-users can refine $f(\cdot)$ based on a preference towards ensuring coverage or small set sizes, we will empirically validate that our methods with either linear scaling (denoted by $\texttt{ECP}_{1}$ / $\texttt{E{\small A}CP}_{1}$) or quadratic scaling (denoted by $\texttt{ECP}_{2}$ / $\texttt{E{\small A}CP}_{2}$) perform well in a wide range of settings.
@@ 4 Experiments
We conduct experiments across a number of large-scale datasets and neural network architectures. Our setup builds on the standard SplitCP procedure introduced in Section 2, which relies on a held-out, in-distribution, “development set” for calibrating the CP threshold. On ImageNet variants, we split the original ImageNet development set (i.e., not used for model training) into a CP calibration set consisting of 25,000 samples, and an in-distribution test set (sometimes called the validation set in the CP literature). The readers are referred to Appendix B for more details.
The conformal threshold is found on the calibration set, and used in subsequent distribution-shifted settings. Importantly, after the conformal threshold is estimated in-distribution, all subsequent steps are unsupervised. We show results on both stationary and continuously shifting test distributions.
@@
We compare our proposed methods to the following baselines:
  • NAIVE: generating prediction sets by including classes until their cumulative softmax score is greater or equal to $1-\alpha$ (the target coverage level). This is generated by the base model itself, without the CP post-processing.
  • Standard SplitCP: applying the CP threshold directly on the distribution-shifted data.
  • SplitCP with ETA: applying the CP threshold while updating the base model using ETA.
Furthermore, in settings with stationary distribution shifts, we compare to Robust Conformal (RC; Cauchois et al. 2024), an existing CP algorithm that handles distribution shifts via robust optimization. In settings with continual distribution shifts, we compare to a number of OCO-based algorithms (Bhatnagar et al., 2023a; Gibbs and Candès, 2024; Zhang et al., 2024) that require additional access to the ground truth labels.
In all experiments, the target coverage rate is set to 0.90. We also analyze our methods with both linear and quadratic scaling, as described in Section 3.3.
Table 1: ECP and EACP can achieve very competitive empirical coverage rates on a number of distribution-shifted datasets, across a variety of imaging domains (ecological, cellular, satellite, etc). All results are from ResNet-50 models except FMOW, which uses a DenseNet-121 (Huang et al., 2016). Quadratic uncertainty scaling provides better coverage rates, however, linear scaling results in smaller set sizes.
Method ImageNet-V2 ImageNet-R ImageNet-A iWildCam RXRX1 FMOW
Coverage SplitCP 0.81 0.50 0.03 0.84 0.84 0.87
NAIVE 0.88 0.69 0.14 0.76 0.48 0.83
RC 0.88 0.63 0.14 0.99 0.91 0.93
ETA 0.81 0.62 0.05 0.84 0.87 0.87
$\texttt{ECP}_{1}$ 0.86 0.61 0.10 0.84 0.87 0.93
$\texttt{ECP}_{2}$ 0.91 0.72 0.27 0.88 0.90 0.96
$\texttt{E{\small A}CP}_{1}$ 0.86 0.71 0.14 0.84 0.90 0.93
$\texttt{E{\small A}CP}_{2}$ 0.91 0.80 0.30 0.89 0.93 0.94
Set Size SplitCP 2.5 3.4 3.4 3.9 81.8 6.2
NAIVE 11.7 20.9 12.7 2.5 6.4 5.8
RC 5.5 10.7 9.6 125 166 10.2
ETA 2.5 3.0 3.6 3.8 100 6.5
$\texttt{ECP}_{1}$ 4.2 9.1 7.4 3.8 105 10.3
$\texttt{ECP}_{2}$ 7.6 23.3 15.1 5.5 137 15.3
$\texttt{E{\small A}CP}_{1}$ 4.5 6.8 8.7 3.7 133 11.1
$\texttt{E{\small A}CP}_{2}$ 8.7 16.1 10.1 5.6 177 16.4
Table 2: Coverage on four different corruption types representing each ImageNet-C category. Compared to the baselines, $\texttt{ECP}_{2}$ closes the coverage gap on most severity levels, while $\texttt{E{\small A}CP}_{2}$ further improves this by achieving the target coverage rate 0.90 on nearly all corruption types and severities.
Method Contrast Brightness Gaussian Noise Motion Blur
1 2 3 4 5 1 2 3 4 5 1 2 3 4 5 1 2 3 4 5
[…]
@@
While the previous datasets present a single distribution shift, the ImageNet-C (Hendrycks and Dietterich, 2019) dataset allows us to investigate shifts across many types and severities. Specifically, ImageNet-C applies 19 visual corruptions to the ImageNet validation set across four corruption categories — noise, blur, weather, and digital, with five severity levels for each corruption. See Appendix B.1 for more details on the datasets.
Stationary shifts
Table 1 summarizes our results on various natural distribution-shifted datasets. We observe that SplitCP (with or without TTA) can exhibit significant gaps with respect to the target coverage rate, whereas ECP closes the gap quite effectively while maintaining meaningful set sizes. Coverage is further improved via EACP, which also helps reducing set sizes on some datasets. In general, we also observe an improvement over RC and NAIVE: the linear scaling variant of our methods has similar coverage rates as these baselines, while the set sizes are typically smaller.
@@
Here we can see the trade-off between linear and quadratic uncertainty scaling. $\texttt{E{\small A}CP}_{2}$ consistently achieves higher coverage rates, however this also leads to “over-coverage” on some datasets and thus larger sets. In contrast, $\texttt{E{\small A}CP}_{1}$ leads to lower coverage but also smaller set sizes. This trade-off can be selected by end-users based on their preference for more accurate or more efficient prediction sets. In subsequent experiments, we will focus on demonstrating the benefit of $\texttt{E{\small A}CP}_{2}$ on coverage, while noting that the observed set sizes are nonetheless practically useful and far from trivial.
@@
In Table 2, we show fine-grained results on one corruption type for each ImageNet-C category, and across each severity level. Here we see the benefit of leveraging an uncertainty notion that can be directly minimized and refined on new test samples. Specifically, $\texttt{E{\small A}CP}_{2}$ is able to recover the target coverage rate on almost all corruption types and severities.